OpenStudy (anonymous):

which is larger 2/3 or 3/4? how do you know? prove it

OpenStudy (anonymous):

3/4 bc when you convert make them have the same denominator it gives you: 2/3=8/12 and 3/4=9/12

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

of course :)

OpenStudy (anonymous):

Check Least common denominator calculator: http://www.hackmath.net/en/calculator/least-common-denominator?input=2%2F3+3%2F4%0D%0A&submit=Calculate+LCD 2/3 = 8/12 3/4 = 9/12 In this case 3 and 4 has no common prime factor that LCD (2/3, 3/4) = 3*4 = 12 so task is reduced to compare numerators 8<9 ;)